OpenStudy (anonymous):

What are the solutions of x^2 + 3x - 18 = 0? Can you please show me the steps? Thanks!

OpenStudy (anonymous):

can you type it for me its hard to read... :d

OpenStudy (anonymous):

haha ok \[x^2 + 3x -18 = (x-3)(x+6) = 0\]

OpenStudy (anonymous):

so x = 3 or x = -6